The magnetic field inside a long solenoid having 50 turns/cm is increased from 2.5 × 10^-3 T to 2.5 T when an iron core of cross-sectional area 4 cm is inserted into it. Find (a) the current in the solenoid, (b) the magnetization I of the core and (c) the pole strength developed in the core.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. B base 1 = 2.5 x 10^-3 B base = 2.5 A = 4 x 10^-4 m^2 , n = 50 turns/cm 5000 turns/m (a) B = μ base 0 ni, ⇒ 2.5 x 10^-3 = 4 π x 10^-1 x 5000 x i ⇒ i = 2.5x 10^-3/ 4π x 10^-7 x 5000 = 0.398 A ≈ 0.4 A (b) I = B base 2 / μ base 0 – H = 2.5 / 4π x 10 ^ -7 – (B base 2 – B base 1) = 2.5 / 4π x 10^-7 – 2.497 = 1.99 x 10 ^ 6 ≈ 2 x 10^6 (c) I = M/V ⇒ ml/Al = m/A ⇒ m = IA = 2 x 10^6 x 4 x 10^-4 = 800 A-m